Your best friend John is willing to sell his brand new car to you at a negotiated price which he is willing to keep unchanged for 15 years. You have paid $5,000 to John to lock the deal. However, you still need to pay $20,000. If you could invest the rest of your savings of $8,000 at an annual interest rate of J1 = 5%, how much money can you accumulate within 15 years and would you be able to buy his car?
1
Expert's answer
2020-06-16T17:32:58-0400

Price of the car is $ 25000

Amount paid is $ 5000

The balance to be paid is $ 20000


"FV=PV(1+r){^n}"


"FV=8000(1+0.05){^1}{^5}"


"FV=8000(1.05){^1}{^5}"


"FV=8000*2.0789"


"FV=16631.43"


The $ 8,000 will have accumulated $ 8,631.43 to a total of $ 16,631.43

Will still fall short of $ 20,000 by $ 3,3368.57 , so would not be able to pay for the car.